OpenStudy (hp34456):

Help!! I'll fan and medal!! What is the arc length if θ = 2 pi over 3 and the radius is 4 cm? 8 pi over 3 16 pi over 3 pi over 6 3 pi over 4

OpenStudy (mathmate):

Arc length = radius * arc-angle (in radians). Example: Arc length with radius 8 and angle pi/2=8*pi/2=4\(\pi\)

OpenStudy (hp34456):

I still don't understand..can you help? @mathmate

OpenStudy (mathmate):

Can you name the are-angle, and the radius?